A 250-kg crate is pushed horizontally with a force of 720 N. If the coefficient of friction is 0.20, calculate the acceleration of the crate.
1
Expert's answer
2012-12-19T11:13:25-0500
F=ma => Newton's law
F=mg - gravity force
F=nmg - friction force
F-ma=nmg
720-250a=0.2*250*10=500
250a=220
a=22/25=0.88 m/s^2